# Long-Context Linear System Identification

### v1

@@ Abstract
This paper addresses the problem of long-context linear system identification, where the state $x_{t}$ of a dynamical system at time $t$ depends linearly on previous states $x_{s}$ over a fixed context window of length $p$. We establish a sample complexity bound that matches the i.i.d. parametric rate up to logarithmic factors for a broad class of systems, extending previous works that considered only first-order dependencies. Our findings reveal a “learning-without-mixing” phenomenon, indicating that learning long-context linear autoregressive models is not hindered by slow mixing properties potentially associated with extended context windows. Additionally, we extend these results to (i) shared low-rank representations, where rank-regularized estimators improve rates with respect to dimensionality, and (ii) misspecified context lengths in strictly stable systems, where shorter contexts offer statistical advantages.
@@ 1 Introduction
System identification, which consists of estimating the parameters of a dynamical system from observations of its trajectories, is a fundamental problem in many fields such as econometrics, robotics, aeronautics, mechanical engineering, or reinforcement learning (Ljung, 1998; Gupta et al., 1976; Moerland et al., 2022). Recent theoretical advances focused on linear system identification, where observations are of the form:
@@
for $t\geqslant 1$, with initialization $x_{0}\in\mathbb{R}^{d}$, noise $\xi_{t}\in\mathbb{R}^{d}$ and design matrix $A^{\star}\in\mathbb{R}^{d\times d}$. Linear system identification (Simpkins, 1999) has been thoroughly studied, with recent interest in sharp non-asymptotic rates (Simchowitz et al., 2018; Sarkar & Rakhlin, 2019; Faradonbeh et al., 2018; Jedra & Proutiere, 2019). The existing analyses, however, focus solely on order-$1$ time dependency, in which the law of $x_{t}$ only depends on the previous state $x_{t-1}$. For order-$p$ time dependencies, the literature on non-asymptotic rates becomes surprisingly scarce, as existing techniques do not extend to $p>1$.
@@
for $t\geqslant p$, the initialization $x_{0},\ldots,x_{p-1}\in\mathbb{R}^{d}$, noise $\xi_{t}\in\mathbb{R}^{d}$ and design matrices $A_{1}^{\star},\ldots,A_{p}^{\star}\in\mathbb{R}^{d\times d}$. This classical $p^{\rm th}$-order vector autoregression model (Box et al., 2015; Brockwell & Davis, 1991; Hamilton, 2020) is termed long-context linear autoregressive model. The term linear refers to the (noisy) linear relationship between iterates and long-context refers to the context length $p$. Recent advances in autoregressive models and architectures such as transformers (Vaswani et al., 2017; Dosovitskiy et al., 2020; El-Nouby et al., 2024) highlight the importance of long-context and its impact on learning. Developing a theoretical understanding of long-context linear autoregressive models is a necessary first step toward tackling these more complex architectures.
@@
Motivated by empirical evidence that high-dimensional data may share some lower-dimensional representation (Bengio et al., 2013; Hospedales et al., 2022), several works additionally studied the problem of learning matrices $A_{k}^{\star}$ under the assumption that they are of low-rank (Alquier et al., 2020; Basu et al., 2019), for order-$1$ autoregressive models. In the long-context setting, this problem is further motivated by the fact that if there exists a lower-dimensional representation of the autoregressive process, this translates into shared kernels for the matrices $A_{k}^{\star}$.
@@
Finally, a key challenge in long-context autoregressive models is misspecification: the system might have an unknown context window $p$ as in Equation 2. $p$ may be arbitrarily large and unknown by the statistician. She may then specify a context length $p^{\prime}$ that can be much smaller, thus yielding the following two fundamental questions: can useful structure still be learned under misspecified context lengths? And what advantages, if any, arise from model misspecification?
Our contributions in long-context linear systems identification are then threefold.
@@
(i) We derive statistical rates on the recovery of matrices $A_{k}^{\star}$ in terms of Frobenius norm, which depends on the number of trajectories $N$ and their length $T$, on the dimension $d$ and the context length $p$. These rates reveal a “learning-without-mixing” phenomenon as they do not have a deflation in effective sample size due to the mixing time of the autoregressive process. This first contribution is an attempt to fill the gap in linear system identification for long context lengths.
@@
(ii) We study statistical guarantees for learning the matrices $A_{k}^{\star}$ assuming that they are all of rank at most $r\ll d$. We prove that the statistical rate reduces, and that rank-regularized estimators adapt to the low-rank structure.
@@
(iii) We study a scenario under which the model is misspecified. Fitting a linear model with context length $p^{\prime}<p$ instead of $p$, we show that the first $p^{\prime}$ matrices are still learned. More importantly, the sample complexity of learning these matrices depends only on the misspecified context length, indicating that misspecification may benefit the model statistically, not just computationally.
Finally, we confirm these statistical rates through experiments that verify the scaling laws predicted by problem parameters. Due to space constraints, these experiments are provided in Appendix C.
@@ 2 Related works
In multivariate linear regression, one observes $\{(x_{i},y_{i})\}_{i=1}^{N}$ from the model $y_{i}=A^{\star}x_{i}+\xi_{i}$, where matrix $A^{\star}\in\mathbb{R}^{d\times d}$ and the sequences of noise $\xi_{i}$ and inputs $x_{i}$ are i.i.d.. The number of samples $N$ needs to scale at least as $d^{2}$ for a good estimation of $A^{\star}$ with ordinary least squares estimator (Hsu et al., 2012; Wainwright, 2019) in Frobenius norm—$\|{A^{\star}-\hat{A}}\|_{F}^{2}\ll 1$. However, in many domains, data is sequential, violating the i.i.d. assumption. In such domains, classical non-i.i.d. formulations, such as vector autoregressive models or discrete-time linear dynamical systems (LDS), as seen in Equation 1, are often employed. Most works used to deal with the non-i.i.d.-ness of the data through mixing time arguments that fall short when the spectral radius of $A^{\star}$ reaches $1$, leading to rates of the form $\|\hat{A}-A^{\star}\|_{F}^{2}=\mathcal{O}(d^{2}/(n(1-\rho)))$ or $\|\hat{A}-A^{\star}\|_{\rm op}^{2}=\mathcal{O}(d/(n(1-\rho)))$ for some spectral quantity $1-\rho$ related to the mixing time of the process. These rates apply to the OLS estimator (Faradonbeh et al., 2018) and online settings (Hardt et al., 2018; Even, 2023) alike.
@@
Simchowitz et al. (2018); Sarkar & Rakhlin (2019) have developed excitation-based arguments to leverage mixing-time independent statistical bounds for the OLS estimator, while Hazan et al. (2017); Jain et al. (2021) respectively used spectral filtering and reverse experience replay in the online setting to obtain such bounds. The estimation of low-rank features has been studied by Basu et al. (2019); Alquier et al. (2020) via nuclear norm regularization. Finally, learning parameters of dynamical systems from $N$ trajectories of length $T$ has previously been considered by Tu et al. (2023) in a more general framework than Equation 1.
@@
Layers of complexity can be added to the LDS described in Equation 1. Mania et al. (2022); Foster et al. (2020) considered non-linear dynamics, that write respectively as $x_{t+1}=A^{\star}\phi(x_{t},u_{t})+\xi_{t}$ and $x_{t+1}=f^{\star}(x_{t})+\xi_{t}$, where in the former $A^{\star}$ is to be estimated and $\phi$ is a known non-linearity, while in the latter $f^{\star}$ is to be estimated. Kostic et al. (2022) recently provided a general framework using Koopman operators, to estimate the parameters of some general Markov chain. Giraud et al. (2015) considered time-varying systems, with arbitrary context lengths, while Bacchiocchi et al. (2024) studied autoregressive bandits. Ziemann & Tu (2022) provided a framework for learning non-parametric dynamical systems with “little mixing”: as their rates are not hindered by slow mixing after a burn-in time (that may itself depend on mixing properties). We refer the reader to Tsiamis et al. (2023) for a survey on recent advances on non-asymptotic system identification of LDS such as in Equation 1. Surprisingly, there does not seem to be much known about long-context LDS in Equation 2, the counterparts of LDS in Equation 1 with a context window $p>1$.
@@ 3 Problem setting
For a matrix $M\in\mathbb{R}^{d_{1}\times d_{2}}$ with singular values $\sigma_{1},\ldots,\sigma_{\min{{\left\{d_{1},d_{2}\right\}}}}$, we denote its squared Frobenius norm as ${{\left\|D\right\|}}_{F}^{2}=\sum_{(i,j)}M_{ij}^{2}=\sum_{\ell}\sigma_{\ell}^{2}$, operator norm as ${{\left\|D\right\|}}_{\rm op}=\max_{\ell}|\sigma_{\ell}|$, and nuclear norm as ${{\left\|D\right\|}}_{*}=\sum_{\ell}|\sigma_{\ell}|$. $I_{d}$ and $0_{d}$ denotes the identity and the null $d\times d$ matrices, respectively. ${\boldsymbol{A}}=\left(A_{1},\ldots,A_{p}\right)$ denotes a rectangular matrix of size $d\times pd$ where each $A_{i}$ is $d\times d$ block.
@@ 3.1 Data generation process
where $x_{s}=0$ for any $s\leqslant 0$ and the noise $\xi_{t}$ is independent of the $x_{s},\xi_{s}$ for $s<t$. This is a particular instance of the general linear autoregressive model in Equation 2 with initial conditions $x_{0},\ldots,x_{p-1}$ set to 00 and the independent noise structure. We assume sub-Gaussian noise:
@@
There exists a known constant $D>0$ such that $\|M_{{\boldsymbol{A}}^{\star}}\|_{\rm op}\leqslant D$.
@@
There is no direct assumption on $L_{\star}$; yet, our results depend on well-behavedness of $\kappa$, the logarithm of the condition number of $L_{\star}$, which is related to $\Gamma_{t}$ that appears in Simchowitz et al. (2018); Sarkar & Rakhlin (2019). $\kappa$ is related to the system stability, as explained in Section 5.
@@
Let $\kappa$ be the logarithm of the condition number of $L_{\star}$, i.e., $\kappa\coloneqq\ln\left(\dfrac{\|L_{\star}\|_{\rm op}}{\sigma_{\min}(L_{\star})}\right)$.
@@ 3.2 Constrained least squares
A natural estimator is the Ordinary Least Square (OLS), defined as any minimizer of the square loss:
@@
is bounded, closed and convex. Hence, the empirical minimizer of the square loss over $\mathcal{A}(D)$ can be computed with projected gradient descent (Duchi et al., 2008) or the Frank-Wolfe algorithm (Jaggi, 2013) as done for $\ell^{1}$ constrained optimization. To avoid projecting onto the set $\mathcal{A}(D)$, following Proposition 3.5, it is possible to restrict $\mathcal{A}(D)$ further into
@@
in order to ensure a condition directly on design matrices. Then, the empirical minimizer of the square loss over $\mathcal{A}(D)^{\prime}$ or $\mathcal{A}(D)^{\prime\prime}$ can again be computed via projected gradient descent or the Frank-Wolfe algorithm, with simplified projection steps.
Lastly, we briefly remark that the diameter constraint in Equation 5 can be removed, i.e., $\mathcal{A}(D)$ replaced by $\mathcal{A}(\infty)$, under an additional assumption on $NT$. This is explained in detail in Section 5.
@@ 3.3 Low-rank assumption
Note that Assumption 3.8 is an instance of Assumption 3.7. The factorization $A_{k}^{\star}=Q^{\star}C_{k}^{\star}$ is another subcase of Assumption 3.7, but is not considered as it leads to iterates that directly lie in the subspace spanned by $Q^{\star}$ and hence $Q^{\star}$ can be learned by treating iterates $x_{t}^{(n)}$ as independent. In order to benefit from the low-rank structure, we consider the following regularized estimator:
@@ 3.4 Misspecification
There exist a constant $D^{\prime}$ such that
@@
Assumption 3.9 is a strong assumption as it requires that $L_{\star}$ is well-behaved regardless of the sequence length $T$. Consequently, the misspecification results are more stringent than other results and apply to a smaller class of systems that still includes strictly stable systems.
@@ 4 Long-context linear system identification
Let Assumptions 3.1 and 3.4 hold. Then, for small $\delta>0$, there exists a constant $C(\delta)=\mathcal{O}(\ln\frac{1}{\delta})$ such that the estimator ${\boldsymbol{\hat{A}}}$ in Equation 5 verifies with probability $1-\delta$:
${{\left\|{\boldsymbol{\hat{A}}}-{\boldsymbol{A}}^{\star}\right\|}}_{F}^{2}\leqslant C(\delta)D^{2}\frac{pd^{2}}{N(T-p)}\mathrm{polylog}(p,d,N,T,\kappa)\,.$ (8)
The explicit constant $C(\delta)$ also depends on the sub-Gaussianity constant $c$ and the diameter $D$ as described in Section B.5 and the rate is numerically verified in Figure 1. Theorem 4.1 exhibits several interesting features.
@@
Second, the rate in Equation 8 exhibits a linear dependency on the context length $p$ instead of a quadratic dependency. This is only due to the number of parameters to be estimated, which is $pd^{2}$ instead of $d^{2}$ and not a deflation in $T$ by a factor of $p$, which implies the context length does not affect the effective sample size. The additive factor in $T-p$ is due to the fact that first iterates do not depend on the full context length, and thus are not as informative as the later iterates. More detailed discussions of Theorem 4.1, in comparison with previous work, can be found in Section 5.
@@
Let Assumptions 3.1, 3.4 and 3.7 hold. Then, for small $\delta>0$, there exists a constant $C(\delta)=\mathcal{O}(\ln\frac{1}{\delta})$ such that the estimator ${\boldsymbol{\hat{A}}}$ in Equation 6 verifies with probability $1-\delta$:
${{\left\|{\boldsymbol{\hat{A}}}-{\boldsymbol{A}}^{\star}\right\|}}_{F}^{2}\leqslant C(\delta)D^{2}\frac{prd}{N(T-p)}\mathrm{polylog}(p,d,r,N,T,\kappa)\,.$ (9)
The improved statistical rate depends on $rd$ instead of $d^{2}$. Note, however, that this estimator cannot be computed in polynomial time, since the underlying optimization problem involves a non-convex constraint on the rank of all $A_{k}$. Several heuristics exist to approximate this estimator. One approach is the Burer-Monteiro factorization (Burer & Monteiro, 2003; 2004), which involves parameterizing $A_{k}$ as $A_{k}=B_{k}C_{k}$ with $B_{k}\in\mathbb{R}^{d\times r}$ and $C_{k}\in\mathbb{R}^{r\times d}$. This method relaxes the constraint to a convex set but results in a non-convex function. Another approach is hard-thresholding algorithms, which use projected (stochastic) gradient descent on the non-convex constraint set (Blumensath & Davies, 2009; Foucart & Subramanian, 2018).
@@
${\boldsymbol{\hat{A}}}\in\operatorname{argmin}{{\left\{\mathcal{L}_{\lambda}({\boldsymbol{A}})\mid{\boldsymbol{A}}\in\mathcal{A}(D)\right\}}}\,,\text{ where}\mathcal{L}_{\lambda}({\boldsymbol{A}})=\mathcal{L}({\boldsymbol{A}})+\lambda{{\left\|{\boldsymbol{A}}\right\|}}_{{*,\rm group}}\,,$ (10)
@@
Let Assumptions 3.1, 3.4, 3.7 and 3.9 hold. Then, for small $\delta>0$, there exists a constant $C(\delta)=\mathcal{O}(\ln\frac{1}{\delta})$ such that the estimator ${\boldsymbol{\hat{A}}}$ in Equation 7 verifies with probability $1-\delta$:
${{\left\|{\boldsymbol{\hat{A}}}-{\boldsymbol{A}}^{\star}_{p^{\prime}}\right\|}}_{F}^{2}\leqslant C(\delta)D^{2}(D^{\prime}+1)^{2}\frac{p^{\prime}dr}{N(T-p)}\mathrm{polylog}(p^{\prime},d,r,N,T,\kappa)\,.$ (11)
For $r=d$, we recover Theorem 4.1 (full-rank setting) for misspecified context windows. The main improvement in that case of Theorem 4.3 over Theorem 4.1 is the dependency on $p^{\prime}$ instead of $p$. In practice, $p$ can be much larger than $p^{\prime}$ and even on the order of $T$. In such a setting, learning all matrices $A_{k}^{\star}$ becomes impossible if $N$ is not large enough and one does not take advantage of the length $T$ of the sequences. One can instead misspecify the student with a context length of $p^{\prime}\ll p$ such that $NT\gg p^{\prime}d^{2}$, so that the first $p^{\prime}$ matrices are still learned.

### v2

@@ Abstract
This paper addresses the problem of long-context linear system identification, where the state $x_{t}$ of a dynamical system at time $t$ depends linearly on previous states $x_{s}$ over a fixed context window of length $p$. We establish a sample complexity bound that matches the i.i.d. parametric rate up to logarithmic factors for a broad class of systems, extending previous works that considered only first-order dependencies. Our findings reveal a “learning-without-mixing” phenomenon, indicating that learning long-context linear autoregressive models is not hindered by slow mixing properties potentially associated with extended context windows. Additionally, we extend these results to (i) shared low-rank representations, where rank-regularized estimators improve the dependence of the rates on the dimensionality, and (ii) misspecified context lengths in strictly stable systems, where shorter contexts offer statistical advantages.
@@ 1 Introduction
System identification, which consists of estimating the parameters of a dynamical system from observations of its trajectories, is a fundamental problem in many fields such as econometrics, robotics, aeronautics, mechanical engineering, and reinforcement learning (Ljung, 1998; Gupta et al., 1976; Moerland et al., 2023). Recent theoretical advances focused on linear system identification, where observations are of the form:
@@
for $t\geqslant 1$, with the initialization $x_{0}\in\mathbb{R}^{d}$, noise $\xi_{t}\in\mathbb{R}^{d}$, and design matrix $A^{\star}\in\mathbb{R}^{d\times d}$. Linear system identification (Simpkins, 1999) has been thoroughly studied, with recent interest in sharp non-asymptotic rates (Simchowitz et al., 2018; Sarkar & Rakhlin, 2019; Faradonbeh et al., 2018; Jedra & Proutière, 2019). The existing analyses, however, focus solely on order-$1$ time dependence, in which the law of $x_{t}$ only depends on the previous state $x_{t-1}$. For order-$p$ time dependencies, the literature on non-asymptotic rates becomes surprisingly scarce, as existing techniques do not extend to $p>1$.
@@
for $t\geqslant p$, the initialization $x_{0},\ldots,x_{p-1}\in\mathbb{R}^{d}$, noise $\xi_{t}\in\mathbb{R}^{d}$ and design matrices $A_{1}^{\star},\ldots,A_{p}^{\star}\in\mathbb{R}^{d\times d}$. This classical $p^{\rm th}$-order vector autoregression model (Box et al., 2015; Brockwell & Davis, 1991; Hamilton, 2020) is termed the long-context linear autoregressive model. The term linear refers to the (noisy) linear relationship between iterates and long-context to the context length $p$. Recent advances in autoregressive models and architectures such as transformers (Vaswani et al., 2017; Dosovitskiy et al., 2021; El-Nouby et al., 2024) highlight the importance of long-context and its impact on learning. Developing a theoretical understanding of long-context linear autoregressive models is a necessary first step toward tackling these more complex architectures.
@@
Motivated by empirical evidence that high-dimensional data may share some lower-dimensional representation (Bengio et al., 2013; Hospedales et al., 2022), several works have also studied the problem of learning matrices $A_{k}^{\star}$ under the assumption that they are low-rank (Alquier et al., 2020; Basu et al., 2019), for order-$1$ autoregressive models. In the long-context setting, this problem is further motivated by the fact that if there exists a lower-dimensional representation of the autoregressive process, this translates into shared kernels for the matrices $A_{k}^{\star}$.
@@
Finally, a key challenge in long-context autoregressive models is misspecification: the system might have an unknown context window $p$ as in Equation 2. $p$ may be arbitrarily large and unknown to the statistician. She may then specify a context length $p^{\prime}$ that can be much smaller, thus yielding the following two fundamental questions: can useful structure still be learned under misspecified context lengths? And what advantages, if any, arise from model misspecification?
Our contributions to long-context linear systems identification are threefold.
@@
(i) We derive statistical rates, which depend on problem parameters $N$, $T$, $p$, and $d$, on the recovery of matrices $A_{k}^{\star}$ in terms of Frobenius norm. These rates reveal a “learning-without-mixing” phenomenon as they do not suffer from a reduction in the effective sample size due to the mixing time of the autoregressive process. This first contribution is an attempt to fill the gap in linear system identification for long context lengths.
@@
(ii) We study statistical guarantees for learning the matrices $A_{k}^{\star}$ assuming that they are all of rank at most $r\ll d$. We prove that the statistical rates improve, and that rank-regularized estimators adapt to the low-rank structure.
@@
(iii) We study a scenario under which the model is misspecified. Fitting a linear model with context length $p^{\prime}<p$ instead of $p$, we show that the first $p^{\prime}$ matrices are still learned. More importantly, the sample complexity of learning these matrices depends only on the misspecified context length, indicating a potential statistical, not just computational, benefit from misspecification. Finally, we confirm these statistical rates through experiments that verify the scaling laws predicted by problem parameters. Due to space constraints, these experiments are provided in Appendix E.
@@ 2 Related works
In multivariate linear regression, one observes $\{(x_{i},y_{i})\}_{i=1}^{N}$ from the model $y_{i}=A^{\star}x_{i}+\xi_{i}$, where the matrix $A^{\star}\in\mathbb{R}^{d\times d}$ and the sequences of noise terms $\xi_{i}$ and inputs $x_{i}$ are i.i.d.. The number of samples $N$ needs to scale at least as $d^{2}$ for a good estimation of $A^{\star}$ with the ordinary least squares estimator (Hsu et al., 2012; Wainwright, 2019) in Frobenius norm—$\|{A^{\star}-\hat{A}}\|_{F}^{2}\ll 1$. However, in many domains, data is sequential, violating the i.i.d. assumption. In such domains, classical non-i.i.d. formulations, such as vector autoregressive models or discrete-time linear dynamical systems (LDS), as seen in Equation 1, are often employed. Traditionally, most works deal with the non-i.i.d.-ness of the data through mixing time arguments that fall short when the spectral radius of $A^{\star}$ reaches $1$, leading to rates of the form $\|\hat{A}-A^{\star}\|_{F}^{2}=\mathcal{O}(d^{2}/(n(1-\rho)))$ or $\|\hat{A}-A^{\star}\|_{\rm op}^{2}=\mathcal{O}(d/(n(1-\rho)))$ for some spectral quantity $1-\rho$ related to the mixing time of the process. These rates apply to the OLS estimator (Faradonbeh et al., 2018) and online settings (Hardt et al., 2018; Even, 2023) alike.
@@
Simchowitz et al. (2018); Sarkar & Rakhlin (2019) develop excitation-based arguments to leverage mixing-time independent statistical bounds for the OLS estimator, while Hazan et al. (2017); Jain et al. (2021) use spectral filtering and reverse experience replay, respectively, in the online setting to obtain such bounds. Basu et al. (2019); Alquier et al. (2020) study the estimation of low-rank features via nuclear norm regularization. Finally, Tu et al. (2024) consider learning parameters of dynamical systems from $N$ trajectories of length $T$ in a more general framework than Equation 1.
@@
Layers of complexity can be added to the LDS described in Equation 1. Mania et al. (2022); Foster et al. (2020) consider non-linear dynamics $x_{t+1}=A^{\star}\phi(x_{t},u_{t})+\xi_{t}$ and $x_{t+1}=f^{\star}(x_{t})+\xi_{t}$, where in the former $A^{\star}$ is to be estimated and $\phi$ is a known non-linearity, while in the latter $f^{\star}$ is to be estimated. Kostic et al. (2022) recently provide a general framework using Koopman operators, to estimate the parameters of some general Markov chain. Giraud et al. (2015) consider time-varying systems, with arbitrary context lengths, while Bacchiocchi et al. (2024) study autoregressive bandits. Ziemann & Tu (2022) provide a framework for learning non-parametric dynamical systems with “little mixing”: as their rates are not hindered by slow mixing after a burn-in time (that may itself depend on mixing properties). We refer the reader to Tsiamis et al. (2023) for a survey of recent advances on non-asymptotic system identification of LDS such as in Equation 1. Surprisingly, there does not seem to be much known about long-context LDS in Equation 2, the counterparts of LDS in Equation 1 with a context window $p>1$.
@@ 3 Problem setting
For a matrix $M\in\mathbb{R}^{d_{1}\times d_{2}}$ with singular values $\sigma_{1},\ldots,\sigma_{\min{{\left\{d_{1},d_{2}\right\}}}}$, we denote its squared Frobenius norm as ${{\left\|M\right\|}}_{F}^{2}=\sum_{(i,j)}M_{ij}^{2}=\sum_{\ell}\sigma_{\ell}^{2}$, operator norm as ${{\left\|M\right\|}}_{\rm op}=\max_{\ell}|\sigma_{\ell}|$, and nuclear norm as ${{\left\|M\right\|}}_{*}=\sum_{\ell}|\sigma_{\ell}|$. $I_{d}$ and $0_{d}$ denote the identity and the null $d\times d$ matrices, respectively. ${\boldsymbol{A}}=\left(A_{1},\ldots,A_{p}\right)$ denotes a rectangular matrix of size $d\times pd$ where each $A_{i}$ is a $d\times d$ block.
@@ 3.1 Data generation process
where $x_{s}=0$ for any $s\leqslant 0$ and the noise $\xi_{t}$ is independent of the $x_{s},\xi_{s}$ for $s<t$. This is a particular instance of the general linear autoregressive model in Equation 2 with initial conditions $x_{0},\ldots,x_{p-1}$ set to 00 and An independent noise structure. We assume sub-Gaussian noise:
@@
There exists a known constant $D\geqslant 1$ such that $\|M_{{\boldsymbol{A}}^{\star}}\|_{\rm op}\leqslant D$.
@@
There is no direct assumption on $L_{\star}$; yet, our results depend on well-behavedness of $\kappa$, the condition number of $L_{\star}$, which is related to $\Gamma_{t}$ that appears in Simchowitz et al. (2018); Sarkar & Rakhlin (2019). $\kappa$ is related to the system stability, as explained in Section 5.
@@
Let $\kappa$ be the condition number of $L_{\star}$, i.e., $\kappa\coloneqq\dfrac{\|L_{\star}\|_{\rm op}}{\sigma_{\min}(L_{\star})}$.
@@ 3.2 Constrained least squares
A natural estimator is the Ordinary Least Squares (OLS), defined as any minimizer of the square loss:
@@
is bounded, closed and convex. Hence, the empirical minimizer of the square loss over $\mathcal{A}(D)$ can be computed with projected gradient descent (Duchi et al., 2008) or the Frank-Wolfe algorithm (Jaggi, 2013) as done for $\ell^{1}$ constrained optimization. To avoid projecting onto the set $\mathcal{A}(D)$, following Proposition 3.1, it is possible to restrict $\mathcal{A}(D)$ further to
@@
in order to ensure a condition directly on the design matrices. Then, the empirical minimizer of the square loss over $\mathcal{A}(D)^{\prime}$ or $\mathcal{A}(D)^{\prime\prime}$ can again be computed via projected gradient descent or the Frank-Wolfe algorithm, with simplified projection steps.
Lastly, we briefly remark that the diameter constraint in Equation 5 can be removed, i.e., $\mathcal{A}(D)$ is replaced by $\mathcal{A}(\infty)$, under an additional assumption on $NT$. This is explained in detail in Section 5.
@@ 3.3 Low-rank assumption
Note that Assumption 3.4 is an instance of Assumption 3.3. The factorization $A_{k}^{\star}=Q^{\star}C_{k}^{\star}$ is another subcase of Assumption 3.3, but is not considered as it leads to iterates that directly lie in the subspace spanned by $Q^{\star}$ and hence $Q^{\star}$ can be learned by treating the iterates $x_{t}^{(n)}$ as independent. In order to benefit from the low-rank structure, we consider the following regularized estimator:
@@ 3.4 Misspecification
There exists a constant $D^{\prime}$ such that
@@
Assumption 3.5 is a strong assumption as it requires that $L_{\star}$ is well-behaved regardless of the sequence length $T$. Consequently, the misspecification results are more restrictive than other results and apply to a smaller class of systems that still includes strictly stable systems as discussed in Section 5.
@@ 4 Long-context linear system identification
Let Assumptions 3.1 and 3.2 hold. Then, for any $0<\delta<e^{-1}$, there exists a constant $C(\delta)$ such that the estimator ${\boldsymbol{\hat{A}}}$ in Equation 5 satisfies with probability $1-\delta$:
${{\left\|{\boldsymbol{\hat{A}}}-{\boldsymbol{A}}^{\star}\right\|}}_{F}^{2}\leqslant C(\delta)D^{2}\frac{pd^{2}}{N(T-p)}\mathrm{polylog}(\kappa,p,d,N,T)\,.$ (8)
The constant $C(\delta)$ depends mildly on the sub-Gaussianity constant $c$ as described in Section C.5 and a sketch of the proof is provided in Appendix A. The rate is numerically verified in Figure 1 in Appendix E. Theorem 4.1 exhibits several interesting features.
@@
Second, the rate in Equation 8 exhibits a linear dependency on the context length $p$ instead of a quadratic dependency. This is only due to the number of parameters to be estimated, which is $pd^{2}$ instead of $d^{2}$ and not a deflation in $T$ by a factor of $p$, which implies the context length does not affect the effective sample size. The additive factor in $T-p$ is due to the fact that the first iterates do not depend on the full context length, and thus are not as informative as the later iterates. More detailed discussions of Theorem 4.1, in comparison with previous work, can be found in Section 5.
@@
Let Assumptions 3.1, 3.2 and 3.3 hold. Then, for any $0<\delta<e^{-1}$, there exists a constant $C(\delta)$ such that the estimator ${\boldsymbol{\hat{A}}}$ in Equation 6 satisfies with probability $1-\delta$:
${{\left\|{\boldsymbol{\hat{A}}}-{\boldsymbol{A}}^{\star}\right\|}}_{F}^{2}\leqslant C(\delta)D^{2}\frac{prd}{N(T-p)}\mathrm{polylog}(\kappa,p,d,r,N,T)\,.$
The improved statistical rate depends on $rd$ instead of $d^{2}$. Note, however, that this estimator cannot be computed in polynomial time, since the underlying optimization problem involves a non-convex constraint on the rank of all $A_{k}$. Several heuristics exist to approximate this estimator. One approach is the Burer-Monteiro factorization (Burer & Monteiro, 2003; 2004), which involves parameterizing $A_{k}$ as $A_{k}=B_{k}C_{k}$ with $B_{k}\in\mathbb{R}^{d\times r}$ and $C_{k}\in\mathbb{R}^{r\times d}$. This method relaxes the constraint to a convex set but results in a non-convex function. Another approach is hard-thresholding algorithms, which use projected (stochastic) gradient descent on the non-convex constraint set (Blumensath & Davies, 2009; Foucart & Subramanian, 2019).
@@
${\boldsymbol{\hat{A}}}\in\operatorname{argmin}{{\left\{\mathcal{L}_{\lambda}({\boldsymbol{A}})\mid{\boldsymbol{A}}\in\mathcal{A}(D)\right\}}}\,,\quad\text{where}\quad\mathcal{L}_{\lambda}({\boldsymbol{A}})=\mathcal{L}({\boldsymbol{A}})+\lambda{{\left\|{\boldsymbol{A}}\right\|}}_{{*,\rm group}}\,,$ (9)
@@
Let Assumptions 3.1, 3.2, 3.3 and 3.5 hold. Then, for any $0<\delta<e^{-1}$, there exists a constant $C(\delta)$ such that the estimator ${\boldsymbol{\hat{A}}}$ in Equation 7 satisfies with probability $1-\delta$:
${{\left\|{\boldsymbol{\hat{A}}}-{\boldsymbol{A}}^{\star}_{p^{\prime}}\right\|}}_{F}^{2}\leqslant C(\delta)D^{2}(D^{\prime}+1)^{2}\frac{p^{\prime}dr}{N(T-p)}\mathrm{polylog}(\kappa,p^{\prime},d,r,N,T)\,.$
For $r=d$, we recover Theorem 4.1 (full-rank setting) for misspecified context windows. In that case, the main improvement of Theorem 4.3 over Theorem 4.1 is the dependency on $p^{\prime}$ instead of $p$. In practice, $p$ can be much larger than $p^{\prime}$ and even on the order of $T$. In such a setting, learning all matrices $A_{k}^{\star}$ becomes impossible if $N$ is not large enough and one does not take advantage of the length $T$ of the sequences. One can instead misspecify the student with a context length of $p^{\prime}\ll p$ such that $NT\gg p^{\prime}d^{2}$, so that the first $p^{\prime}$ matrices are still learned.